8V, 3AH BATTERY OPERATING A DEVICE FOR 5.5HRS. WHAT IS THE ENERGY USED IN JOULES?
Answer #1

Energy of battery in joules = q*V

where q is charge and given by current(i)*time(sec)

V is voltage (volts)

Energy of battery in joules = 3*5.5*3600*8 = 475.20 kJ